OpenStudy (anonymous):

How do I write these questions, 5^square root a^2 and 1/5 square root of x^4 using exponents?

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt[5]{a^2}= (a^2)^1/5 = a^2/5\]

OpenStudy (anonymous):

note: in the above 1/5 is a power, and 2/5 is a power (even though may not look like this).

OpenStudy (anonymous):

\[1/\sqrt[5]{x^4}= 1/ ((x^4)^1/5 ) = (x^4)^ -1/5 = x^-4/5\]

OpenStudy (anonymous):

note again: 1/5, -1/5 and -4/5 are all powers even though may not look like it!